Phone Home, Inc. is considering a new 4-year expansion project that requires an initial fixed asset investment of $3 million. The fixed asset will be depreciated straight-line to zero over its 4-year tax life, after which time it will have a market value of $225,000. The project requires an initial investment in net working capital of $330,000, all of which will be recovered at the end of the project. The project is estimated to generate $2,640,000 in annual sales, with costs of $1,056,000. The tax rate is 33 percent and the required return for the project is 15 percent.

If you do not have enough cash to purchase the fixed asset for the project, but want to maintain a debt-to-equity ratio of . 5. It will cost 8% to raise equity and 6% to raise debt. If you take into account flotation costs for fixed asset, what is the npv of the project.
Answer #1

Calculating cash outflow in year 0

Initial investment in fixed asset = total capital = $3 million = 3 x 1000000 = $3000000

Debt / equity = 0.5

Therefore we get debt/Total capital = 0.5/1.5 = 1/3

Equity/Total capital = 1 - (debt/Total capital ) = 1 - (1/3) = 2/3

Total debt raised = Total capital x (debt / total capital) = 3000000 x (1/3) = 1000000

Total equity raised = Total capital x (equity / total capital) = 3000000 x (2/3) = 2000000

As flotation are tax deductible, therefore

Net flotation cost for debt = cost of raising debt x total debt raised (1- tax rate) = 6% x 1000000 (1-33%) = 6% x 1000000 x 67% = 40200

Net Flotation cost for equity = Cost of raising equity x total equity raised (1-tax rate) = 8% x 2000000 (1 - 33%) = 8% x 2000000 x 67% = 107200

Initial investment in working capital = 330000

Cash outflow in year 0 = - Initial investment in fixed asset - Net flotation cost for debt - Net flotation cost for equity - investment in working capital = -3000000 - 40200 - 107200 - 330000 = -3477400

Calculating After tax Operating cash flow

annual straight line depreciation = (Initial investment in fixed asset - Book value at end of 4 years) / No of years = (3000000 - 0) / 4 = 3000000 / 4 = 750000

After tax Operating cash flow for a year 1 to 4 = EBIT(1-tax rate) + depreciation = (Sales - costs - depreciation)(1-tax rate) + depreciation

= (2640000 - 1056000 - 750000)(1-33%) + 750000 = 834000 x 67% + 750000 = 558780 + 750000 = 1308780

Calculating Terminal cash flow at end of year 4

Book value of equipment at end of 4 years = 0

Salvage value of equipment at end of 4 years = 225000

Terminal cash flow in year 4 = Salvage value of equipment at end of 4 years + Recovery of investment in working capital - tax on gain from sale of fixed asset

= Salvage value of equipment at end of 4 years + Recovery of investment in working capital - tax rate(Salvage value of equipment at end of 4 years - book value of equipment at end of 4 years)

= 225000 + 330000 - 33%(225000 - 0) = 225000 + 330000 - 74250 = 480750

Calculating NPV

Year 1 2 3 4
After tax operating cash flow (a) 1308780 1308780 1308780 1308780
Terminal cash flow (b) 480750
Net Cash flow = (a) + (b) 1308780 1308780 1308780 1789530

Net present value of project = Initial cash outflow in year 0 + Sum of present value of Net cash flow for year 1 to 4 discounted at required rate of return of the project

= -3477400 + 1308780 / (1 + 15%) + 1308780 / (1 + 15%)2 + 1308780 / (1 + 15%)3 + 1789530 / (1 + 15%)4

= -3477400 + 1138069.5652 + 989625.7088 + 860544.0946 + 1023169.5855

= 534008.9541 = 534008.95

Hence NPV of the project taking into account flotation cost of asset = 534008.95